# Line and surface defects in 5D $N=2$ SCFT from matter-coupled $F(4)$ gauged supergravity

### v2

@@ 1 Introduction
In this work, we generalize these two classes of solutions to the matter-coupled $F(4)$ gauged supergravity. We will consider two gauged supergravities obtained from coupling pure or minimal $F(4)$ gauged supergravity to three and one vector multiplets. The former leads to $SO(4)\sim SO(3)_{R}\times SO(3)$ gauge group admitting two supersymmetric $AdS_{6}$ vacua with $SO(3)_{R}\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries [25], see also [26, 27, 28, 29] for other hologrpahic solutions obtained from this gauged supergravity. $SO(3)_{R}$ denotes the R-symmetry while the second $SO(3)$ factor arises from the symmetry of the three vector multiplets. On the other hand, the latter gauged supergravity leads to $SO(3)_{R}\times U(1)$ gauge group. The $U(1)$ factor is just the abelian gauge symmetry of the single vector multiplet under which all the other fields are uncharged. A number of supersymmetric Janus solutions and $AdS_{6}$ black holes from this gauged supergravity has been studied more recently in [30, 31, 32]. For similar holographic solutions describing conformal surface defects in seven-dimensional gauged supergravities, see [33, 34, 35, 36].
@@ 2 Matter coupled $N=(1,1)$ gauged supergravity in six dimensions
$e^{-1}\mathcal{L}$ $=$ $\frac{1}{4}R-e\partial_{\mu}\sigma\partial^{\mu}\sigma-\frac{1}{4}P^{I\alpha}_{\mu}P^{\mu}_{I\alpha}-\frac{1}{8}e^{-2\sigma}\mathcal{N}_{\Lambda\Sigma}\widehat{F}^{\Lambda}_{\mu\nu}\widehat{F}^{\Sigma\mu\nu}-\frac{3}{64}e^{4\sigma}H_{\mu\nu\rho}H^{\mu\nu\rho}$
@@
3 Line defects from matter-coupled $F(4)$ gauged supergravtiy
In this section, we consider solutions of the matter-coupled $F(4)$ gauged supergravity which are holographically dual to conformal line defects within $N=2$ SCFTs in five dimensions. The metric ansatz can be written as
@@ 3 Line defects from matter-coupled $F(4)$ gauged supergravtiy
$\gamma^{\alpha}=\rho^{\alpha}\otimes\sigma_{2}\mathbb{I}_{2},\qquad\gamma^{r}=\rho_{*}\otimes\sigma_{2}\otimes\mathbb{I}_{2},$
@@ 3.2 Line defects with $SO(2)$ symmetry
$F_{(3,2)}$ $=$ $-\frac{288c_{6}^{2}|dk|}{g_{1}^{2}}\left\{\frac{\rho|dk|}{\bar{\Delta}}b(\tilde{r})\wedge\left[(\sinh^{2}\phi_{2}-\cosh^{2}\phi_{2}\cos^{2}\tilde{\theta})\textrm{vol}_{\tilde{S}^{2}}\right.\right.$
@@
4 Surface defects from matter-coupled $F(4)$ gauged supergravtiy
We now move to another type of solutions with the metric ansatz given by
@@ 4 Surface defects from matter-coupled $F(4)$ gauged supergravtiy
$\gamma^{\alpha}=\rho^{\alpha}\otimes\mathbb{I}_{2}\otimes\sigma,\qquad\gamma^{r}=-(\mathbb{I}_{2}\otimes\gamma_{*}\otimes\sigma_{1}),$
@@ 5 Conclusions
It would be interesting to identify line and surface defects, in terms of position dependent deformations, dual to the solutions found here in the framework of five-dimensional SCFTs. It is also interesting to use holographic renormalization to compute the expectation values of the defect operators and boundary stress tensor as done in [20] for the case of line defect solution found in [18] and match with the field theory results. Finally, finding type IIB solutions describing possible brane configurations, of which the uplifted solutions given here describe low-energy regimes as in the case of pure $F(4)$ gauged supergravity studied in [18] and [19], is also worth considering. We note that although the six-dimensional solutions for both line and surface defects are the same, the uplifted solutions could correspond to different brane configurations in type IIB theory. In particular, the uplifted solution for line defects would involve F1-D1-branes, due to the presence of two-form fluxes along $AdS_{2}$, and D3-branes corresponding to a non-trivial four-form field on top of the 5-brane web dual to the SCFT in five dimensions [40]. On the other hand, we expect the brane configuration corresponding to the surface defects to involve only D3-branes without F1-D1-branes since the electric part of the two-form field does not appear in the uplifted ten-dimensional solution.
@@ Appendix B Consistent truncation of type IIB theory on $S^{2}\times\Sigma$
In this appendix, we collect relevant formulae for uplifting solutions of six-dimensional $F(4)$ gauged supergravity coupled to one vector multiplet to type IIB theory. In particular, we mainly consider $SO(2)_{R}$ singlet sector of the $F(4)$ gauged supergravity with two scalars denoted by $\phi_{0}$ and $\phi_{2}$ and the two-form field. Although the solutions found in this paper only involve non-vanishing $\phi_{2}$, we give all the formulae of the truncation for both $\phi_{0}$ and $\phi_{2}$ non-vanishing. The result can be useful in other applications such as uplifting the Janus solution found in [30] by setting the two-form field to zero.
@@
$C^{a}_{(2)}$ $=$ $-\frac{24c_{6}}{g_{1}^{2}}\textrm{vol}_{\tilde{S}^{2}}(k^{a}+L^{a})-\frac{24c_{6}}{g_{1}^{2}}\frac{|dk|^{2}}{\bar{\Delta}}\cosh\phi_{0}\cosh\phi_{2}\sin^{2}\tilde{\theta}\sigma^{a}\wedge d\tilde{\varphi}-4c_{6}p^{a}B,$

### v4

@@ 1 Introduction
In this work, we generalize these two classes of solutions to the matter-coupled $F(4)$ gauged supergravity. We will consider two gauged supergravities obtained from coupling pure or minimal $F(4)$ gauged supergravity to three and one vector multiplets. The former leads to $SO(4)\sim SO(3)_{R}\times SO(3)$ gauge group admitting two supersymmetric $AdS_{6}$ vacua with $SO(3)_{R}\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries [25], see also [26, 27, 28, 29] for other hologrpahic solutions obtained from this gauged supergravity. $SO(3)_{R}$ denotes the R-symmetry while the second $SO(3)$ factor arises from the symmetry of the three vector multiplets. On the other hand, the latter gauged supergravity leads to $SO(3)_{R}\times U(1)$ gauge group. The $U(1)$ factor is just the abelian gauge symmetry of the single vector multiplet under which all the other fields are uncharged. A number of supersymmetric Janus solutions and $AdS_{6}$ black holes from this gauged supergravity has been studied more recently in [30, 31, 32]. For similar holographic solutions describing conformal surface defects in seven-dimensional gauged supergravities, see [33, 34, 35, 36, 37].
@@ 2 Matter coupled $N=(1,1)$ gauged supergravity in six dimensions
$e^{-1}\mathcal{L}$ $=$ $\frac{1}{4}R-\partial_{\mu}\sigma\partial^{\mu}\sigma-\frac{1}{4}P^{I\alpha}_{\mu}P^{\mu}_{I\alpha}-\frac{1}{8}e^{-2\sigma}\mathcal{N}_{\Lambda\Sigma}\widehat{F}^{\Lambda}_{\mu\nu}\widehat{F}^{\Sigma\mu\nu}-\frac{3}{64}e^{4\sigma}H_{\mu\nu\rho}H^{\mu\nu\rho}$
@@
3 Line defects from matter-coupled $F(4)$ gauged supergravity
In this section, we consider solutions of the matter-coupled $F(4)$ gauged supergravity which are holographically dual to conformal line defects within $N=2$ SCFTs in five dimensions. The procedure to find these solutions closely follows that given in [18]. The metric ansatz can be written as
@@ 3 Line defects from matter-coupled $F(4)$ gauged supergravity
$\gamma^{\alpha}=\rho^{\alpha}\otimes\sigma_{2}\otimes\mathbb{I}_{2},\qquad\gamma^{r}=\rho_{*}\otimes\sigma_{2}\otimes\mathbb{I}_{2},$
@@ 3.2 Line defects with $SO(2)$ symmetry
$F_{(3,2)}$ $=$ $-\frac{288c_{6}^{2}|dk|}{g_{1}^{2}}\left\{\frac{\rho|dk|}{\bar{\Delta}}b^{\prime}(\tilde{r})d\tilde{r}\wedge\textrm{vol}_{AdS_{2}}\wedge\left[(\sinh^{2}\phi_{2}-\cosh^{2}\phi_{2}\cos^{2}\tilde{\theta})\textrm{vol}_{\tilde{S}^{2}}\right.\right.$
@@
4 Surface defects from matter-coupled $F(4)$ gauged supergravity
We now move to another type of solutions describing two-dimensional conformal defects within five-dimensional $N=2$ SCFTs. Following [19], we use the metric ansatz of the form
@@ 4 Surface defects from matter-coupled $F(4)$ gauged supergravity
$\gamma^{\alpha}=\rho^{\alpha}\otimes\mathbb{I}_{2}\otimes\sigma_{2},\qquad\gamma^{r}=-(\mathbb{I}_{2}\otimes\gamma_{*}\otimes\sigma_{1}),$
@@ 5 Conclusions
It would be interesting to identify line and surface defects, in terms of position dependent deformations, dual to the solutions found here in the framework of five-dimensional SCFTs. It is also interesting to use holographic renormalization to compute the expectation values of the defect operators and boundary stress tensor as done in [20] for the case of line defect solution found in [18] and match with the field theory results. Finally, finding type IIB solutions describing possible brane configurations, of which the uplifted solutions given here describe low-energy regimes as in the case of pure $F(4)$ gauged supergravity studied in [18] and [19], is also worth considering. We note that although the six-dimensional solutions for both line and surface defects are the same, the uplifted solutions could correspond to different brane configurations in type IIB theory. In particular, the uplifted solution for line defects would involve F1-D1-branes, due to the presence of two-form fluxes along $AdS_{2}$, and D3-branes corresponding to a non-trivial four-form field on top of the 5-brane web dual to the SCFT in five dimensions [41, 42]. On the other hand, we expect the brane configuration corresponding to the surface defects to involve only D3-branes without F1-D1-branes since the electric part of the two-form field does not appear in the uplifted ten-dimensional solution.
@@ Appendix B Consistent truncation of type IIB theory on $S^{2}\times\Sigma$
In this appendix, we collect relevant formulae for uplifting solutions of six-dimensional $F(4)$ gauged supergravity coupled to one vector multiplet to type IIB theory as given in [38]. In particular, we mainly consider $SO(2)_{R}$ singlet sector of the $F(4)$ gauged supergravity with two scalars denoted by $\phi_{0}$ and $\phi_{2}$ and the two-form field. Although the solutions found in this paper only involve non-vanishing $\phi_{2}$, we give all the formulae of the truncation for both $\phi_{0}$ and $\phi_{2}$ non-vanishing. The result can be useful in other applications such as uplifting the Janus solution found in [30] by setting the two-form field to zero.
@@
$C^{a}_{(2)}$ $=$ $-\frac{24c_{6}}{g_{1}^{2}}\textrm{vol}_{\tilde{S}^{2}}(k^{a}+L^{a})$
$-\frac{24c_{6}}{g_{1}^{2}}\frac{|dk|^{2}}{\bar{\Delta}}\cosh\phi_{0}\cosh\phi_{2}\sin^{2}\tilde{\theta}\sigma^{a}\wedge d\tilde{\varphi}-4c_{6}p^{a}B,$